OpenStudy (anonymous):

Finish these three solutions to this equation. 3x-5y=-3 (__, 5), (1, __), (0, __)

OpenStudy (anonymous):

1.) x= (-3+25)/3=22/3 2.)y=(3+3)/5=6/5 3.)y=3/5